OpenStudy (anonymous):

what is the solution to the equation log3*x - log3*2 = 2 a. x = 18 b. x = 12 c. x = 9/2 d. x = 4/3

zepdrix (zepdrix):

\[\large \log_3x-\log_32=2\]Using rules of logarithms allows us to combine the logs like so, \[\large \log_3\left(\frac{x}{2}\right)=2\]From here we can rewrite it in exponential form, let me know if this step is confusing.\[\large 3^2=\frac{x}{2}\] From here, do you understand how to solve for x? :)

OpenStudy (anonymous):

@zepdrix im sorry I don't :/

zepdrix (zepdrix):

Hmm it looks like the left side gives us 9 since 3 squared is 3x3. \[\large 9=\frac{x}{2}\] Multiplying both sides by 2 will put it in the form x= which is what we want :)

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hmm multiplying by 2 gives us,\[\large 9\cdot2=\frac{x}{\cancel2}\cdot \cancel2\] Which simplifies to,\[\large 18=x\]

OpenStudy (anonymous):

Oh okay! thanks a lot! I really appreciate the help :)